# Intermediate Mass Ratio Inspirals in Milky Way Galaxies

### v1

@@ 1 Introduction
LISA is a gravitational wave detector which will launch in the mid-2030s and operate for 4-10 years. Its 2.5 million km baseline makes it able to detect mergers of black holes in the intermediate mass range to redshift $z=20$ and beyond, including phenomena such as Milky Way-like SMBHs consuming intermediate mass black holes (IMBHs) delivered by merging dwarf galaxies. This type of merger is likely to have a mass ratio of $\lesssim 0.01$, which is known as an Intermediate Mass Ratio Inspiral, or IMRI (defined as $10^{-2}>q>10^{-5}$) (Amaro-Seoane et al., 2023). More specifically it is known as a “heavy IMRI,” which is an SMBH - IMBH merger. Gravitational wave detections of IMRIs have the unique capability to determine black hole masses within a few percent; observations such as these will allow tight constraints on dwarf MBH masses and possibly provide limits for seed formation mechanisms as well. However, progress is needed in waveform modeling to allow for the proper interpretation of gravitational wave signals from these IMRI events. See section 3.2 of Amaro-Seoane et al. (2023) for more details on IMRIs of all types.
@@ 2.2 Black Hole Physics
where $\rho(<{\bf v}_{\rm BH})$ is the density of collisionless particles moving slower than the BH, $M_{\rm BH}$ and $v_{\rm BH}$ are the mass and velocity of the BH, respectively, and ln $\Lambda$ is the Coulomb logarithm. This latter quantity depends on the impact parameters $b_{min}$ and $b_{max}$, such that ln $\Lambda\sim$ ln ($\frac{b_{max}}{b_{min}}$). We set $b_{max}$ equal to the softening length, because dynamical friction is well-resolved on scales this size and larger. For the minimum impact parameter, we set it to be the minimum deflection radius, $b_{min}=GM_{BH}/v_{\rm BH}^{2}$, with a minimum possible value of the Schwarzschild Radius. The resulting acceleration is added to the BH’s current acceleration and integrated at every timestep, such that dynamical effects from structures smaller than our resolution element can be approximately accounted for. We note that while the BH within a galaxy is subject to this dynamical friction model, dyamical friction BH’s host dwarf galaxy inspiralling into a larger host is captured by the simulation and modeled correctly via the $N$-body method.
Our simulations do not resolve the full inspiral process of a BH pair; instead, we merge BH particles together when the equivalent of a close pair has formed. BH particles merge when they come within less than two softening lengths and also meet the criterion $\frac{1}{2}\Delta{\vec{\rm v}}<\Delta{\vec{\rm a}}\cdot\Delta{\vec{\rm r}}$, where $\Delta{\vec{\rm v}},\Delta{\vec{\rm a}}$ and $\Delta{\vec{\rm r}}$ represent the relative velocity, acceleration, and radius vectors of the two BHs respectively. This latter criterion ensures that the BHs have low relative velocities and prevent mergers during e.g. rapid flybys. When the BHs merge, their masses are summed and the remnant is positioned at the center of mass of the system.
@@ 2.3 Limitations of the Model
Another limitation of our model, which is due to resolution, is that the BH merger methodology described above omits a great deal of physics. In our model, the final few $\sim 100$ pc of simulated binary coalescence occurs instantaneously, but in reality several processes govern the actual timescale. For example, dynamical friction due to background matter acts on scales of 1 - 100 pc. At closer separations, loss-cone scattering of stars in the central region shrinks the separation of black holes, as does emission of gravitational radiation. The presence of gas can also affect the coalescence of binary black holes at all scales. Each of these processes has its own timescale on which it acts, depending on the environment of the binary, but none of them are so quick to make the timescale negligible. Estimates of the binary hardening time-scale in massive galaxies range from $10^{7}-10^{8}$ years (Armitage & Natarajan, 2002; Haiman et al., 2009; Colpi, 2014; Holley-Bockelmann & Khan, 2015), though full coalescence timescales from large distances can be much longer depending on mass ratio and total system mass (Katz et al., 2020; Volonteri et al., 2020). This timescale can become even longer in less massive galaxies, due to the nature of the shallower potential wells of dwarf galaxies. De Cun et al. (2023) showed that BH mergers in dwarf galaxies can take up to a few $10^{9}$ years due to inefficient dynamical friction in galaxies with cored density profiles (though some estimates for low-mass galaxies are shorter (Khan & Holley-Bockelmann, 2021) and some longer (Tamfal et al., 2018)). In this work we present the inspiral time of mergers of IMBHs with SMBHs throughout cosmic time, often in low-mass galaxies, but our simulations do not take the abovementioned delays into account. We will discuss the repercussions of this issue in the following sections.
@@ 3 Demographics of the Mergers
The distribution of black hole merger mass ratios is quite skewed towards small ($<0.1$) values. Figure 1 shows a histogram of the log of the mass ratios, which has an approximately bimodal appearance. There is a prevalence of ratios with values $-2.2<\log q<1.5$ or 1:150 $<q<$1:30, indicating a predominance of mergers in the IMRI regime. The remainder of the mergers span the 1:10 $<q<$ 1:1 space, representing more equal-mass mergers. We use this bimodality to delineate IMRIs from more equal-mass mergers, and henceforth color IMRIs as yellow in subsequent figures, and non-IMRIs as purple. While the break in the distribution occurs at a 1:25 ratio or $q\sim 0.04$, rather than $q<0.01$ as defined by the community, we argue that this definition is reasonable because (a) our mass ratios are uncertain to a factor of a few (Section 2.3) and (b) the computational challenges of simulating IMRI waveforms extend to all $q$ values less than 0.1 (Section 5).
Figure 2: Distribution of mergers times in Gyr since the big bang. The purple histogram is all black hole mergers, while overplotted in yellow are the IMRI mergers. The IMRI subset follows approximately the same distribution as the total.
@@
The majority of the mergers take place within the early universe – about 3 Gyr after the Big Bang. Figure 2 shows the distribution of merger times, exhibiting a large peak at the earliest times. The purple histogram shows all black hole mergers, while overplotted in yellow are the IMRI mergers (identified from the left peak of Figure 1).These times do not account for unresolved delays in BH coalescence (see Section 2.3). We expect these delays should be on the order of a maximum of 1 Gyr for close to equal-mass mergers in low mass galaxies (De Cun et al., 2023), but it is possible it could be longer for IMRI-type mergers.
@@
Figure 3: Distribution of inclinations (far left), eccentricities (center left), angular momentum (center right), and merger durations in Gyr (far right). The purple histograms represent all mergers, while yellow denote IMRIs. This color scheme is used throughout the paper.
@@
The initial orbital parameters of each MBH is dependent on how its host galaxy enters the main halo, and this entry governs the overall evolution of much of the inspiral and merger. To investigate these details, we measure the initial inclination, angular momentum, and eccentricity of the entry of each MBH into the main halo. The time of halo entry is determined by the first simulation snapshot where the BH is within the virial radius of the main halo. At this moment the host dwarf galaxy may still be identified as an independent galaxy, before its tidal disruption is complete. The resolution of this process is dependent on the time resolution of the snapshots, which is on average about 50 Myr. Because the earliest BH mergers happen very quickly in small galaxies, their halo entry and inspiral are not resolved and a full analysis is not possible. Therefore, we are forced to omit some of the earliest mergers in our subsequent results for lack of time resolution.
@@
We also measure the initial eccentricity of BHs as they enter the halo, shown in the center right panel of Figure 3. We calculate eccentricity by tracing the orbit of each BH after it crosses the virial radius of the main halo, and measuring its pericenter and apocenter distances. Some inspiral examples are shown in Figure 4, which shows the infall of an example black hole over time. We calculate eccentricity using the following formula:
@@
The primary quantity which correlates with inspiral time is the compactness of the host galaxy.We devise a compactness parameter as a general proxy for density, and calculate it by finding the half-mass radius of the stellar population of each host galaxy, and dividing the stellar mass within that radius by the cube of that radius at the time of infall. At this moment the dwarf host has already undergone considerable stripping, and the virial radius closely resembles the stellar radius of the galaxy. In the top panel of Figure LABEL:fig:durations, we show that more compact galaxies have shorter infall times compared to more diffuse ones, as one expects due to the nature of dynamical friction. A more compact galaxy can plunge deeper into a larger halo before being disrupted, resulting in a MBH which is closer to the center and will inspiral more quickly.
@@
Other orbital parameters show no apparent trends with inspiral time. For example, the duration of inspiral does not seem to have a strong dependence on the initial eccentricity or angular momentum of the BH orbit, with both circular and radial orbits having a variety of inspiral times. The same holds true for the mass ratio of the merger, most of which take 2-4 Gyr to inspiral and merge regardless of the value of $q$. And while compactness plays a role, the independent properties of dwarf galaxy stellar mass and galaxy radius are not correlated with inspiral time in any way on their own.
@@ 4 Evolution of the IMRIs
The inspirals in the top two panels of 6 are good examples of an initially highly eccentric orbit becoming more circular as it evolves. This decrease in eccentricity is also seen in Weller et al. (2023). However some orbits exhibit more chaotic behavior and do not appear to circularize - examples are shown in the two bottom panels of the figure. Even though the pericenter decreases over time, the eccentricity does not decrease. In Figure 7 we show the distribution of the change in eccentricities, with negative values becoming more circular. About half of the inspirals become more circular with time, while the remainder are unchanged by the dynamical influences of the inspiral.
@@ 5 LISA repercussions
In this work we have shown that $\sim$50of MBH mergers in Milky Way environments constitute IMRIs. Mergers of these demographics should be detectable by the LISA observatory, but are less straightforward to characterize compared to major mergers (0.1 $<q<$1.0) or extreme mass ratio inspirals (EMRIs; $q<10^{-5}$). For gravitational wave signals to be efficiently identified, they are compared to an existing library of modeled waveforms. Major merger waveforms are fairly straightforward to model because one can use the Post-Newtonian (PN) approximation of the numerical relativity calculation to acquire waveforms (Blanchet, 2014). For EMRIs, defined as $10^{-8}<q<10^{-5}$, waveform models are approachable because the small mass ratio lends itself to the application of a perturbative approach based on a systemic expansion of the field equations (Barack & Pound, 2019). However, IMRIs straddle these two regimes and neither approach is fully applicable. Possibly one can use a combination, using perturbative and PN formalisms in domains where each are valid. Full numerical relativity solutions are extremely computationally intensive due to the extremely high temporal and spatial resolutions required, and are difficult to compute for the vast array of IMRI parameter space (Duez & Zlochower, 2019).
@@
The importance of highlighting this problem cannot be overstated. Since a large fraction of mergers in complex galaxy environments may be IMRIs, the LISA community must develop the necessary tools to detect and analyze such signals. It is likely that new approaches must be developed to generate the variety of waveforms needed to fully analyze the LISA data when it comes online (e.g. Rink et al., 2024). The black hole masses, redshifts, and frequencies of these ‘heavy’ IMRI events can reveal much about the formation of high-redshift SMBH seeds, as well as the assembly of massive galaxies and their SMBHs. With these signals we can determine much about the formation of the building blocks of our universe, but not if we cannot interpret LISA’s observations. An organized, concerted effort is needed to develop waveform codes and further predict the demographics of IMRI populations, so that by the mid 2030’s our community will be ready for LISA’s breakthroughs.

### v2

@@ 1 Introduction
LISA is a gravitational wave detector which will launch in the mid-2030s and operate for 4-10 years (Amaro-Seoane et al., 2017; Colpi et al., 2024). Its 2.5 million km baseline makes it able to detect mergers of black holes in the intermediate mass range to redshift $z=20$ and beyond, including phenomena such as Milky Way-like SMBHs consuming intermediate mass black holes (IMBHs) delivered by merging dwarf galaxies. This type of merger is likely to have a mass ratio of $\lesssim 0.01$, which is known as an Intermediate Mass Ratio Inspiral, or IMRI (defined as $10^{-2}>q>10^{-5}$) (Amaro-Seoane et al., 2023). More specifically it is known as a “heavy IMRI,” which is an SMBH - IMBH merger. Gravitational wave detections of IMRIs have the unique capability to determine black hole masses within a few percent; observations such as these will allow tight constraints on dwarf MBH masses and possibly provide limits for seed formation mechanisms as well. However, progress is needed in waveform modeling to allow for the proper interpretation of gravitational wave signals from these IMRI events. See section 3.2 of Amaro-Seoane et al. (2023) for more details on IMRIs of all types.
@@ 2.2 Black Hole Physics
where $\rho(<{\bf v}_{\rm BH})$ is the density of collisionless particles moving slower than the BH, $M_{\rm BH}$ and $v_{\rm BH}$ are the mass and velocity of the BH, respectively, and ln $\Lambda$ is the Coulomb logarithm. This latter quantity depends on the impact parameters $b_{min}$ and $b_{max}$, such that ln $\Lambda\sim$ ln ($\frac{b_{max}}{b_{min}}$). We set $b_{max}$ equal to the softening length, because dynamical friction is well-resolved on scales this size and larger. For the minimum impact parameter, we set it to be the minimum deflection radius, $b_{min}=GM_{BH}/v_{\rm BH}^{2}$, with a minimum possible value of the Schwarzschild Radius. The resulting acceleration is added to the BH’s current acceleration and integrated at every timestep, such that dynamical effects from structures smaller than our resolution element can be approximately accounted for. We note that while the BH within a galaxy is subject to this dynamical friction model, dynamical friction BH’s host dwarf galaxy inspiralling into a larger host is captured by the simulation and modeled correctly via the $N$-body method.
Our simulations do not resolve the full inspiral process of a BH pair; instead, we merge BH particles together when the equivalent of a close pair has formed. BH particles merge when they come within less than two softening lengths and also meet the criterion $\frac{1}{2}\Delta{\vec{\rm v}}<\Delta{\vec{\rm a}}\cdot\Delta{\vec{\rm r}}$, where $\Delta{\vec{\rm v}},\Delta{\vec{\rm a}}$ and $\Delta{\vec{\rm r}}$ represent the relative velocity, acceleration, and radius vectors of the two BHs respectively. This latter criterion ensures that the BHs have low relative velocities and prevent mergers during e.g. rapid flybys. When the BHs merge, their masses are summed and the remnant is positioned at the center of mass of the system. This merger methodology is a crude and fairly inaccurate representation of the actual physics of black hole mergers, which include a gravitational recoil kick as well as mass loss in the system from radiated gravitational wave energy. Recoil kicks can range from $10-1000$ km s-1 which is often enough to eject a remnant black hole out of its host galaxy at high redshifts, when escape velocities are small. The effect of gravitational recoil in cosmological simulations has been studied in a post-processing capacity by Dunn et al. (2020) and on-the-fly by Dong-Páez et al. (2024), and both find that massive black hole growth is stifled when black holes are ejected from galaxies and no longer contribute to the black hole mass function. However, for IMRI events with low mass ratios, both the mass loss and the recoil kick are predicted to be smaller (Campanelli et al., 2007; Baker et al., 2008). IMRI recoil kicks of $\sim 10$km s-1 would have little effect on disturbing the remnant’s position from the center of the host galaxy.
@@ 2.3 Limitations of the Model
Another limitation of our model, which is due to resolution, is that the BH merger methodology described above omits a great deal of physics. In our model, the final few $\sim 100$ pc of simulated binary coalescence occurs instantaneously, but in reality several processes govern the actual timescale. For example, dynamical friction due to background matter acts on scales of 1 - 100 pc. At closer separations, loss-cone scattering of stars in the central region shrinks the separation of black holes, as does emission of gravitational radiation. The presence of gas can also affect the coalescence of binary black holes at all scales. Each of these processes has its own timescale on which it acts, depending on the environment of the binary, but none of them are so quick to make the timescale negligible. To highlight the importance of this multitude of processes, in their absence it would take $\sim 10^{30}$ years for gravitational emission alone to bring black holes from the scale of our simulation resolution to merger (see Chan et al. (2018) equation 4), underscoring the need for further simulations to resolve and document these timescales more accurately.
Estimates of the binary hardening time-scale in massive galaxies range from $10^{7}-10^{8}$ years (Armitage & Natarajan, 2002; Haiman et al., 2009; Colpi, 2014; Holley-Bockelmann & Khan, 2015), though full coalescence timescales from large distances can be much longer depending on mass ratio and total system mass (Katz et al., 2020; Volonteri et al., 2020). This timescale can become even longer in less massive galaxies, due to the nature of the shallower potential wells of dwarf galaxies. De Cun et al. (2023) showed that BH mergers in dwarf galaxies can take up to a few $10^{9}$ years due to inefficient dynamical friction in galaxies with cored density profiles (though some estimates for low-mass galaxies are shorter (Khan & Holley-Bockelmann, 2021) and some longer (Tamfal et al., 2018)). In this work we present the inspiral time of mergers of IMBHs with SMBHs throughout cosmic time, often in low-mass galaxies, but our simulations do not take the abovementioned delays into account. We will discuss the repercussions of this issue in the following sections.
@@ 3 Demographics of the Mergers
The distribution of black hole merger mass ratios is quite skewed towards small ($<0.1$) values. Figure 1 shows a histogram of the log of the mass ratios, which has an approximately bimodal appearance. There is a prevalence of ratios with values $-2.2<\log q<-1.5$ or 1:150 $<q<$1:30, indicating a predominance of mergers in the IMRI regime. The remainder of the mergers span the 1:10 $<q<$ 1:1 space, representing more equal-mass mergers. We use this bimodality to delineate IMRIs from more equal-mass mergers, and henceforth color IMRIs as yellow in subsequent figures. While the break in the distribution occurs at a 1:25 ratio or $q\sim 0.04$, rather than $q<0.01$ as defined by the community, we argue that this definition is reasonable because (a) our mass ratios are uncertain to a factor of a few (Section 2.3) and (b) the computational challenges of simulating IMRI waveforms extend to all $q$ values less than 0.1 (Section 5).
Figure 2: Log of mass ratio vs merger times in Gyr since the big bang. The purple circles represent non-IMRI mergers, while yellow stars are the IMRI mergers. Both subsets show similar activity with time, with increased mergers at the earliest epochs.
@@
The majority of the mergers take place within the early universe – about 3 Gyr after the Big Bang. Figure 2 shows the log of the mass ratio vs time of merger, exhibiting the highest merger rates at the earliest times. The purple circles show all non-IMRI black hole mergers, while the yellow stars are the IMRI mergers (identified from the left peak of Figure 1). These times do not account for unresolved delays in BH coalescence (see Section 2.3). We expect these delays should be on the order of a maximum of 1 Gyr for close to equal-mass mergers in low mass galaxies (De Cun et al., 2023), but it is possible it could be longer for IMRI-type mergers.
@@
Figure 3: Distribution of inclinations (far left), eccentricities (center left), angular momentum (center right), and merger durations in Gyr (far right). The purple histograms represent all mergers, while histograms overplotted in yellow denote IMRIs. This color scheme is used throughout the paper.
@@
The initial orbital parameters of each MBH is dependent on how its host galaxy enters the main halo, and this entry governs the overall evolution of much of the orbital inspiral and merger. To investigate these details, we measure the initial inclination, angular momentum, and eccentricity of the entry of each MBH into the main halo. The time of halo entry is determined by the first simulation snapshot where the BH is within the virial radius of the main halo. At this moment the host dwarf galaxy may still be identified as an independent galaxy, before its tidal disruption is complete. The resolution of this process is dependent on the time resolution of the snapshots, which is on average about 50 Myr. Because the earliest BH mergers happen very quickly in small galaxies, their halo entry and inspiral are not resolved and a full analysis is not possible. Therefore, we are forced to omit some of the earliest mergers in our subsequent results for lack of time resolution.
@@
We also measure the initial eccentricity of BHs as they enter the halo, shown in the center right panel of Figure 3. We calculate eccentricity by tracing the orbit of each BH after it crosses the virial radius of the main halo, and measuring its pericenter and apocenter distances. Some orbital inspiral examples are shown in Figure 4, which shows the infall of an example black hole over time. We calculate eccentricity using the following formula:
@@
The primary quantity which correlates with inspiral time is the compactness of the host galaxy.We devise a compactness parameter as a general proxy for density, and calculate it by finding the half-mass radius of the stellar population of each host galaxy, and dividing the stellar mass within that radius by the cube of that radius at the time of infall. At this moment the dwarf host has already undergone considerable stripping, and the virial radius closely resembles the stellar radius of the galaxy. In the top panel of Figure 5, we show that more compact galaxies have shorter infall times compared to more diffuse ones, as one expects due to the nature of dynamical friction. A more compact galaxy can plunge deeper into a larger halo before being disrupted, resulting in a MBH which is closer to the center and will inspiral more quickly.
@@
Other orbital parameters show no apparent trends with orbital inspiral time. For example, the duration of inspiral does not seem to have a strong dependence on the initial eccentricity or angular momentum of the BH orbit, with both circular and radial orbits having a variety of inspiral times. The same holds true for the mass ratio of the merger, most of which take 2-4 Gyr to inspiral and merge regardless of the value of $q$. And while compactness plays a role, the independent properties of dwarf galaxy stellar mass and galaxy radius are not correlated with inspiral time in any way on their own.
@@ 4 Evolution of the IMRIs
The inspirals in the top two panels of 6 are good examples of an initially highly eccentric orbit becoming more circular as it evolves. This decrease in eccentricity is also seen in Weller et al. (2023). However some orbits exhibit more chaotic behavior and do not appear to circularize - examples are shown in the two bottom panels of the figure. Even though the pericenter decreases over time, the eccentricity does not decrease. In Figure 7 we show the distribution of the change in eccentricities, with negative values becoming more circular. About half of the inspirals become more circular with time due to galactic-scale dynamical processes, while the remainder are unchanged by the dynamical influences of the inspiral. Gravitational wave emission can be an additional circularizing effect, however, which is not accounted for here (Peters, 1964; Hinder et al., 2008); our predictions should be taken as lower limits on eccentricity evolution.
@@ 5 LISA repercussions
In this work we have shown that $\sim$50of MBH mergers in Milky Way environments constitute IMRIs. Mergers of these demographics should be detectable by the LISA observatory, but are less straightforward to characterize compared to major mergers (0.1 $<q<$1.0) or extreme mass ratio inspirals (EMRIs; $q<10^{-5}$). For gravitational wave signals to be efficiently identified, they are compared to an existing library of modeled waveforms. Major merger waveforms are fairly straightforward to model because one can use the Post-Newtonian (PN) approximation of the numerical relativity calculation to acquire waveforms (Blanchet, 2014). For EMRIs, defined as $10^{-8}<q<10^{-5}$, waveform models are approachable because the small mass ratio lends itself to the application of a perturbative approach based on a systemic expansion of the field equations (Barack & Pound, 2019). However, IMRIs straddle these two regimes and neither approach is fully applicable. Possibly one can use a combination, using perturbative and PN formalisms in domains where each are valid.
Full numerical relativity solutions are extremely computationally intensive due to the extremely high temporal and spatial resolutions required, and are difficult to compute for the vast array of IMRI parameter space (Duez & Zlochower, 2019; LISA Consortium Waveform Working Group et al., 2023). Numerical methods have been used to directly calculate waveforms for mass ratios up to $q=0.001$, but in the case of head-on black hole collisions, not quasi-circular inspirals Lousto & Healy (2023). These simulations are a good proof of concept, but are not yet astrophysically relevant. In the case of quasi-circular orbits, numerical relativity simulations have attained waveforms for mass ratios of 1:15 to 1:20 (Jani et al., 2016; Boyle et al., 2019), but for fewer cycles than are need for a LISA inspiral, and lacking higher-order modes. In the event that a LISA-detected IMRI has high signal-to-noise and exhibits higher modes, there are no models of sufficient resolution to capture them (Ferguson et al., 2021). In terms of hybrid approaches, some methods may use approaches such as post-adiabatic corrections or small mass-ratio perturbation theory to reach intermediate mass ratios from the EMRI modeling regime. However, these estimates do not include the actual merger aspect of the waveform (e.g. van de Meent & Pfeiffer, 2020; Katz et al., 2021) and are therefore not useful for analyzing actual IMRI merger events.
The importance of highlighting this problem cannot be overstated. While the nature of the zoom-in simulations we study here does not allow for a direct IMRI rate calculation, we can generally state that Milky Way-type galaxies are likely to host a handful of IMRI events throughout their evolution. While $\sim$ half of these may occur at redshifts $z>5$ and have too low signal-to-noise to be detected (see Figure 2), the remainder are likely LISA sources. Larger mass galaxies with more satellites may exhibit even more IMRIs. Since a large fraction of mergers in complex galaxy environments may be IMRIs, the LISA community must develop the necessary tools to detect and analyze such signals. It is likely that new approaches must be developed to generate the variety of waveforms needed to fully analyze the LISA data when it comes online (e.g. Rink et al., 2024). IMRIs will add complexity to the overall signal and must be included in the global fit to properly interpret all LISA data.
@@
The black hole masses, redshifts, and frequencies of these ‘heavy’ IMRI events can reveal much about the formation of high-redshift SMBH seeds, as well as the assembly of massive galaxies and their SMBHs. With these signals we can determine much about the formation of the building blocks of our universe, but not if we cannot interpret LISA’s observations. An organized, concerted effort is needed to develop waveform codes and further predict the demographics of IMRI populations, so that by the mid 2030’s our community will be ready for LISA’s breakthroughs.